# LLMs Homogenize Values in Constructive Arguments on Value-Laden Topics

### v1

@@ Abstract.
Large language models (LLMs) are increasingly used to promote prosocial and constructive discourse online. Yet little is known about how they negotiate and shape underlying values when reframing people’s arguments on value-laden topics. We conducted experiments with 347 participants from India and the United States, who wrote constructive comments on homophobic and Islamophobic threads, and reviewed human-written and LLM-rewritten versions of these comments. Our analysis shows that LLM systematically diminishes Conservative values while elevating prosocial values such as Benevolence and Universalism. When these comments were read by others, participants opposing same-sex marriage or Islam found human-written comments more aligned with their values, whereas those supportive of these communities found LLM-rewritten versions more aligned with their values. These findings suggest that LLM-driven value homogenization can shape how diverse viewpoints are represented in contentious debates on value-laden topics and may influence the dynamics of online discourse critically.
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Human-centered computing Empirical studies in HCI
@@ 1. Introduction
Although values strongly shape how people frame their arguments on divisive topics (Ajjour et al., 2019; Chen et al., 2019), little is known about how LLMs handle underlying values when rewriting people’s comments constructively. Given the growing adoption of LLMs in providing writing support (Zhang et al., 2023, 2025a; Biswas et al., 2025; Siddiqui et al., 2025) and facilitating civil discourse online (Johnson, 2024; Meta, 2024; LinkedIn, 2024), it is essential to understand how these models affect the expression of values in value-laden topics. If LLM rewrites comments in ways that shift or distort underlying socio-cultural values, its output may appear constructive but risks alienating the very people it aims to support. Therefore, in this work we aim to explore:
@@
    What values do people from different cultures emphasize when writing constructively on value-laden topics?
@@
    How might human–LLM value differences affect which comments people prefer?
To find answer, we conducted a three-phase study with 347 participants from India and the United States, recruited via Prolific. In Phase 1, participants first shared their perceptions of constructiveness and wrote constructive comments on the same homophobic and Islamophobic threads. In Phase 2, we introduced the intervention: we used GPT-4 to rewrite participants’ comments from Phase 1 constructively. We also invited additional participants, who wrote an initial constructive comment and then used GPT-4 for rewriting their comments constructively. Together, these steps yielded our pairs of human-written and LLM-rewritten constructive comments. We analyzed the underlying values in these comments based on Schwartz’s theory of basic human values (Schwartz, 1994). Finally, in Phase 3 we evaluated perceived alignment. A new group of participants, distinct from those in Phases 1 and 2, reviewed the human-written and LLM-rewritten comment pairs and indicated which better aligned with their own values.
In response to RQ1, we found that people’s stances on homophobia, rather than culture, influenced what values they expressed when writing constructively on homophobic threads. Participants opposing same-sex marriage framed their arguments through Traditional values more (see Figure 1) than those who supported. Following RQ2, when an LLM was employed to rewrite comments constructively, it systematically suppressed Conservative values while amplifying prosocial values, such as Universalism and Benevolence (i.e., welfare of others and the society; see Figure 1). In this process, LLM changed the stance towards more neutral and supportive direction, especially for comments that were written by opponents of same-sex marriage and Islam. As a result, in RQ3, when these comments were reviewed by other people, those opposing same-sex marriage and Islam found human-written comments more aligned with their values, while those supportive favored LLM-rewritten versions. We discuss the political tensions and ethical implications of LLM-driven value homogenization when shaping online discourse on value-laden topics. Taken together, our work makes the following contributions:
@@
  • We point to the ethical issue of LLMs distorting stances by rewriting comments from individuals who oppose same-sex marriage or Islam as if they were supportive, thereby misrepresenting and conflicting with their value systems.
@@ 2. Related Work
We situate our work by first describing research on using LLMs to facilitate constructive arguments on value-laden topics, outlining the role of values in shaping argumentation. We then present existing work examining human-LLM value alignments.
@@ 2.1. Constructive Disagreements and The Role of Values
Constructive disagreements are typically characterized as civil dialogues that are focused, issue-relevant (Sukumaran et al., 2011; Kolhatkar and Taboada, 2017b) and evidence-based (Kolhatkar and Taboada, 2017a). While other definitions focus on outcomes, such as dispute resolution (De Kock and Vlachos, 2021), finding common ground (Friess and Eilders, 2015), improved team performance (Niculae and Danescu-Niculescu-Mizil, 2016), or the emergence of new solutions (Friess and Eilders, 2015). In online discourse, constructiveness manifests through specific linguistic features, such as longer content (Sukumaran et al., 2011; Kolhatkar and Taboada, 2017b), richer argumentative markers (e.g., discourse connectives, stance adverbials, reasoning verbs, modals, and root clauses) (Kolhatkar and Taboada, 2017a; Napoles et al., 2017), greater readability (Kolhatkar et al., 2020), greater use of named entities (Kolhatkar et al., 2020; Niculae and Danescu-Niculescu-Mizil, 2016), more politeness (Napoles et al., 2017; Zhang et al., 2018; De Kock and Vlachos, 2021), and less hedging (Niculae and Danescu-Niculescu-Mizil, 2016; Napoles et al., 2017; De Kock et al., 2022).
Given the tension, time, and effort involved in expressing constructive disagreement on value-laden topics (Cutler et al., 2022; Mun et al., 2024; Gurgun et al., 2023), researchers have explored if LLMs can support people during this process. For example, Shahid et al. (2025) found that LLMs can help people express their opinions on value-laden topics more constructively than they could do on their own, by integrating different linguistic markers of constructiveness in writing. Similarly, Argyle et al. (2023) demonstrated that LLMs can help people reframe their arguments by incorporating politeness, restating opposing views, or validating others’ sentiments, which improve conversation quality during disagreement. Kambhatla et al. (2024) also showed that reframing human-written comments by an LLM to include hedging, acknowledgment, and agreement enabled readers to be more open to alternative perspectives on value-laden topics. Additionally, Govers et al. (2024) observed that highly cooperative and persuasive strategies present in LLM-mediated comments were effective in influencing reader’s opinions on polarizing value-laden topics. This growing body of work highlights the promise of LLMs in scaffolding constructive disagreement by leveraging linguistic strategies that individuals often struggle to apply in contentious situations.
@@ 3. Methods
To address our research questions, we conducted a three-phase study (see Figure 2) with participants from two different cultures: India and the United States (US), who are known to differ in value priorities (Singh et al., 1962). In the first phase, we examined which values participants emphasize when writing constructively on value-laden topics. In the second phase, we analyzed how these underlying values change when LLM rewrites people’s comments constructively. In the third phase, we investigated which type of comment: human or LLM-written people find more aligned with their values. The study protocol was approved by the institutional review board at our institution.
3.1. Phase 1: Values in Human-Written Constructive Comments
@@
Study Design. We recruited 58 Indian and 58 American participants on Prolific. Each participant was asked to write constructive comments on value-laden topics. For this writing task, we curated two Reddit threads: one related to homophobia (Those who oppose same-sex marriage, why?) and another related to Islamophobia (Why do people dislike Islam?). We chose these issues because they are highly value-laden and considered socially divisive both in India and the US. Each thread, presented in English, included the original post and four user comments. Overall, 63% of comments in these threads were negative in sentiment (average: -0.69), with several being highly negative and toxic.
@@
We first asked participants about their stance on same-sex marriage and Islam, using adapted instruments from Pew research survey (Borelli, 2022; Center, 2024) since people’s values are associated with their stances on divisive issues (Mason, 2015). We then randomly showed participants either a homophobic or an Islamophobic thread, and asked them to write a constructive comment in response to the thread. To ensure comparability, we used the same threads for both demographics, so that any differences in values could not be attributed to variation in stimuli. We also asked participants an open-ended question about their perceptions of constructive comments to get more insight into how they framed their writing. We compensated each participant with USD 1.00 for completing the writing task.
@@
Data Analysis. We collected 116 constructive comments written by our participants. We first used GPT-4 to annotate the values expressed in different segments of each comment, following Schwartz’s theory of basic human values (Schwartz, 1994)—which has been used to detect values in arguments in cross-cultural setting (Kiesel et al., 2022). Additionally, several researchers applied Schwartz’s theory to measure human-LLM value alignment (Shen et al., 2024, 2025; Liu et al., 2025; Wang et al., 2025; Segerer, 2025). Schwartz’s theory identifies a set of universal values that guide human behavior, organized into ten broad categories such as Security, Tradition, Conformity, Universalism, Benevolence, Power, Achievement, Stimulation, Hedonism, and Self-direction (Schwartz, 1994).
@@
We used GPT-4 to annotate these values in participants’ comments. Overall, each comment represented a vector of values. For example, the comment shown in Figure 3 represents six distinct values: Universalism, Hedonism, Self-direction, Tradition, Conformity, and Security. The annotations can either have a positive or negative association with the value. For example, the segment ‘why society should interfere’ captures lack of tolerance (Universalism). Some segments can represent multiple values, such as ‘man get married to a woman’ signifies both Traditional value and Conformity to rules.
@@
Two authors independently reviewed GPT-4’s annotation of values, making updates when we disagreed with the assigned value or when a comment segment reflected multiple values. Inter-rater reliability was high (Cohen’s $\kappa=0.87$). Any remaining disagreements were resolved through discussion. After the annotations were confirmed, we conducted multiple Chi-square tests with Bonferroni correction to examine if there was any cross-cultural difference in the values Indian and Americans expressed in their comments across different topics and stances.
@@ 3.2. Phase 2: Values in LLM-Rewritten Constructive Comments
LLM Rewriting. Next, we used GPT-4 to rewrite participants’ comments from Phase 1 to make them constructive. To minimize misalignment between human’s and LLM’s understanding of constructiveness, we relied on the different characteristics of constructive comments that participants provided in Phase 1. We synthesized their open-ended responses into a summary definition: constructive comments are “respectful, fact-based, balanced, and thoughtful contributions that promote understanding and progress in conversation rather than conflict”–which also align with the perceptions of constructiveness from prior studies (Sukumaran et al., 2011; Friess and Eilders, 2015; Kolhatkar and Taboada, 2017a; Napoles et al., 2017; Zhang et al., 2018). We used zero-shot, cultural prompting to ensure that LLM rewrites algin with the corresponding country’s cultural values (Tao et al., 2024). We instructed GPT-4 to rewrite each comment constructively within 100 words, matching the average length of constructive comments as reported in prior studies (Kolhatkar and Taboada, 2017b). We used the following prompt:
@@
For each human-written comment from Phase 1, we generated one LLM-rewritten version, resulting in 116 LLM-rewritten constructive comments. Since we used a fixed prompt to generate these rewrites, we wanted to make sure that our findings are not sensitive to the prompt. Therefore, we collected additional comments from participants, where they could choose their own prompts to rewrite constructively using LLM, which we describe below.
Human-LLM Collaborative Writing. Similar to Phase 1, we recruited 51 new participants (US: 26, India: 25) on Prolific for writing constructive comments on value-laden topics. Participants first reported their stance on same-sex marriage and Islam since people’s values and stances on value-laden topics are closely related (Mason, 2015). Then they were asked to write a constructive comment in response to a randomly selected homophobic and an Islamophobic thread. We used different sets of threads (see Table 7 in Appendix) from Phase 1 to ensure that our analyses on value expressions were not sensitive to specific threads.
@@
We asked participants to write a constructive comment on their own first. Then they could prompt an LLM to rewrite their comments constructively (see Figure 4). We provided a set of prompts based on different characteristics of constructive comments that other participants described in Phase 1. Participants could select multiple prompts or write their own prompt. We used GPT-4 to rewrite their comments constructively in real-time using the following instruction:
@@
Data Analysis. We generated 116 LLM-rewritten version of the comments that participants wrote in Phase 1. Additionally, from human-LLM collaborative writing in Phase 2, we collected 102 human-written comments along with LLM-rewritten versions. Overall, this led to 218 pairs of human-written comments and their corresponding LLM-rewritten versions.
First, we wanted to verify if the LLM-rewritten versions were indeed more constructive compared to the original human-written comments. Following prior work (Shahid et al., 2025; Kolhatkar and Taboada, 2017b; Napoles et al., 2017; Zhang et al., 2018), we calculated eight linguistic features of constructiveness, including length, readability score, politeness, number of named entities, and argumentative features, such as number of discourse connectives, stance adverbials, reasoning verbs, modals, and root clauses. For length, we calculated the number of words in each comment. We used SMOG index (Park et al., 2023) (textstat library in Python) to measure readability scores. We evaluated politeness scores using politeness package in R. We utilized SpaCy (Honnibal et al., 2017) to get number of named entities and root clauses. For discourse connectives, stance adverbials, reasoning verbs, and modals, we adapted the open-source code from Kolhatkar et al. (2020). After calculating the features, we performed Mann-Whitney tests with Bonferroni correction to inspect how these features differed between original human-written and LLM-rewritten comments. Finally, to examine how LLMs handle underlying values when rewriting constructively, we annotated the values in our collected comments based on Schwartz’s theory of basic human values (Schwartz, 1994). We followed the same procedure as described above in Phase 1. We performed multiple CHI-square tests with Bonferroni correction to analyze whether the values differed between human-written and LLM-rewritten constructive comments.

### v2

@@ Abstract.
Large language models (LLMs) are increasingly used to promote prosocial and constructive discourse online. Yet little is known about how these models negotiate and shape underlying values when reframing people’s arguments on value-laden topics. We conducted experiments with 465 participants from India and the United States, who wrote comments on homophobic and Islamophobic threads, and reviewed human-written and LLM-rewritten constructive versions of these comments. Our analysis shows that LLM systematically diminishes Conservative values while elevating prosocial values such as Benevolence and Universalism. When these comments were read by others, participants opposing same-sex marriage or Islam found human-written comments more aligned with their values, whereas those supportive of these communities found LLM-rewritten versions more aligned with their values. These findings suggest that value homogenization in LLM-mediated prosocial discourse runs the risk of marginalizing conservative viewpoints on value-laden topics and may inadvertently shape the dynamics of online discourse.
@@
journalyear: 2026copyright: ccconference: Proceedings of the 2026 CHI Conference on Human Factors in Computing Systems; April 13–17, 2026; Barcelona, Spainbooktitle: Proceedings of the 2026 CHI Conference on Human Factors in Computing Systems (CHI ’26), April 13–17, 2026, Barcelona, Spaindoi: 10.1145/3772318.3791624isbn: 979-8-4007-2278-3/2026/04ccs: Human-centered computing Empirical studies in HCI
@@ 1. Introduction
Although values greatly shape how people frame their arguments on divisive topics (Ajjour et al., 2019; Chen et al., 2019), little is known about how LLMs handle underlying values when rewriting people’s comments constructively. Given the growing adoption of LLMs in providing writing support (Zhang et al., 2023, 2025a; Biswas et al., 2025; Siddiqui et al., 2025) and facilitating civil discourse online (Johnson, 2024; Meta, 2024; LinkedIn, 2024), it is essential to understand how these models affect the expression of values in value-laden topics. If LLM rewrites comments in ways that shift or distort underlying socio-cultural values, its output may appear constructive but risks alienating the very people it aims to support. Therefore, in this work we explore:
@@
    What values do people from different cultures emphasize when writing comments on value-laden topics?
@@
    How might differences between human–LLM value framings affect which comments people prefer?
To find answer, we conducted a three-phase study with 465 participants from India and the United States, recruited via Prolific. In Phase 1, participants responded to the same homophobic and Islamophobic threads by writing either a regular comment (baseline condition) or a constructive comment (constructive condition). In Phase 2, we introduced the intervention: we used GPT-4 to rewrite participants’ comments from Phase 1 constructively. We also invited additional participants, who wrote an initial comment and then used GPT-4 for rewriting their comments constructively. Together, these steps yielded our pairs of human-written and LLM-rewritten constructive comments. We analyzed the underlying values in these comments based on Schwartz’s theory of basic human values (Schwartz, 1994). Finally, in Phase 3 we evaluated perceived alignment. A new group of participants, distinct from those in Phases 1 and 2, reviewed the human-written and LLM-rewritten comment pairs and indicated which one better aligned with their own values.
In response to RQ1, we found that people’s stances on homophobia, rather than demographic factors, such as culture, age, gender, or religion, influenced what values they expressed when commenting on homophobic threads. Participants opposing same-sex marriage framed their arguments by emphasizing Traditional values more (see Figure 1) than those who supported. Following RQ2, when an LLM was employed to rewrite comments constructively, it systematically suppressed Conservative values while amplifying prosocial values, such as Universalism and Benevolence (i.e., welfare of others and the society; see Figure 1). In this process, LLM frequently changed the stance towards more neutral and supportive direction, especially for comments that were written by opponents of same-sex marriage and Islam. As a result, in RQ3, when these comments were reviewed by different people, those opposing same-sex marriage and Islam found human-written comments more aligned with their personal values, while those supportive favored LLM-rewritten versions. We discuss the political tensions and ethical implications of LLM-driven value homogenization when shaping online discourse on value-laden topics. Taken together, our work makes the following contributions:
@@
  • We point to the ethical tension when LLM distorts stances by rewriting comments from individuals who oppose same-sex marriage or Islam as if they were supportive, thereby misrepresenting and conflicting with their value systems.
@@ 2. Related Work
We situate our work first by describing research on using LLMs to facilitate constructive arguments on value-laden topics, outlining the role of values in shaping argumentation. We then present existing work examining human-LLM value alignments.
@@ 2.1. Constructive Disagreements and The Role of Values
Constructive disagreements are typically characterized as civil dialogues that are focused, issue-relevant (Sukumaran et al., 2011; Kolhatkar and Taboada, 2017b) and evidence-based (Kolhatkar and Taboada, 2017a). Other definitions focus on conversational outcomes, such as dispute resolution (De Kock and Vlachos, 2021), finding common ground (Friess and Eilders, 2015), improved team performance (Niculae and Danescu-Niculescu-Mizil, 2016), or the emergence of new solutions (Friess and Eilders, 2015). In online discourse, constructiveness manifests through specific linguistic features, such as longer content (Sukumaran et al., 2011; Kolhatkar and Taboada, 2017b), richer argumentative markers (e.g., discourse connectives, stance adverbials, reasoning verbs, modals, and root clauses) (Kolhatkar and Taboada, 2017a; Napoles et al., 2017), greater readability (Kolhatkar et al., 2020), greater use of named entities (Kolhatkar et al., 2020; Niculae and Danescu-Niculescu-Mizil, 2016), more politeness (Napoles et al., 2017; Zhang et al., 2018; De Kock and Vlachos, 2021), and less hedging (Niculae and Danescu-Niculescu-Mizil, 2016; Napoles et al., 2017; De Kock et al., 2022).
Given the tension, time, and effort involved in expressing constructive disagreement on value-laden topics (Cutler et al., 2022; Mun et al., 2024; Gurgun et al., 2023), researchers have explored if LLMs can support people during this process. For example, Shahid et al. (2025) found that LLMs can help people express their opinions on value-laden topics more constructively than they could do on their own, by integrating different linguistic markers of constructiveness in writing. Similarly, Argyle et al. (2023) demonstrated that LLMs can help people reframe their arguments by incorporating politeness, restating opposing views, or validating others’ sentiments, which improve conversation quality during disagreement. Kambhatla et al. (2024) also showed that reframing human-written comments by an LLM to include hedging, acknowledgment, and agreement–enabled readers to be more open to alternative perspectives on value-laden topics. Additionally, Govers et al. (2024) observed that highly cooperative and persuasive strategies present in LLM-mediated comments were effective in influencing reader’s opinions on polarizing value-laden topics. This growing body of work highlights the promise of LLMs in scaffolding constructive disagreement by leveraging linguistic strategies that individuals often struggle to apply in contentious situations.
@@ 3. Methods
To address our research questions, we conducted a three-phase study (see Figure 2) with participants from two different cultures: India and the United States (US). We selected these countries as representatives of individualist (US) and collectivist (India) cultures, who are known to have different value priorities (Singh et al., 1962). In the first phase, we examined which values participants emphasize when writing comments on value-laden topics. In the second phase, we analyzed how these underlying values change when LLM rewrites people’s comments constructively. In the third phase, we investigated what type of comment: human or LLM-written people find more aligned with their values. The study protocol was approved by the institutional review board at our institution.
3.1. Phase 1: Values in Human-Written Comments
@@
Study Design. We recruited 117 Indian and 117 American participants on Prolific. Participants were randomly assigned to one of two conditions: a baseline condition or a constructive condition. In the baseline condition, participants were asked to write a comment on a value-laden topic. In the constructive condition, they were explicitly instructed to write a constructive comment. For this writing task, we curated two Reddit threads: one related to homophobia (Those who oppose same-sex marriage, why?) and another related to Islamophobia (Why do people dislike Islam?). We chose homophobia and Islamophobia because these issues are highly value-laden and socially divisive in both India and the US. Apart from their cross-cultural relevance, research shows that discussions involving anti-Muslim and anti-LGBTQ sentiment tend to elicit greater toxicity than non-identity related controversial topics, such as climate change and vaccination (Salminen et al., 2020; Seckin et al., 2025), making these issues suitable for our study.
We first asked participants about their demographics, such as age, gender, and religion they identify with. We then asked their stance on same-sex marriage and Islam, using adapted instruments from Pew research survey (Borelli, 2022; Center, 2024) since people’s values are associated with their stances on divisive issues (Mason, 2015). Next, participants were randomly shown either a homophobic or an Islamophobic thread. Each thread included the original post in English and four user comments. We manually selected these comments from Reddit threads that captured diverse stances while maintaining a toxic tone, reflecting realistic online hostility. Overall, 63% of these comments were negative in sentiment (average: -0.69), with several being highly negative and toxic.
@@
Depending on the condition, participants either wrote a comment (baseline) or a constructive comment (constructive) in response to the thread. To ensure comparability, we used the same threads for participants in both conditions so that any differences in values could not be attributed to variation in stimuli. We also asked participants in the constructive condition an open-ended question about their perceptions of constructive comments to get more insight into how they framed their writing. We compensated each participant with USD 1.00 for completing the writing task.
Evaluating Constructiveness. We collected 234 comments (118 baseline and 116 constructive) written by our participants. To assess whether prompting to “write constructively” led to any difference, we compared the linguistic markers of constructiveness in comments from the baseline and constructive conditions. Following prior work (Shahid et al., 2025; Kolhatkar and Taboada, 2017b; Napoles et al., 2017; Zhang et al., 2018), we calculated eight measures of constructiveness, including length, readability score, politeness, number of named entities, and argumentative features, such as number of discourse connectives, stance adverbials, reasoning verbs, modals, and root clauses. For length, we calculated the number of words in each comment. We used SMOG index (Park et al., 2023) (textstat library in Python) to measure readability scores. We evaluated politeness scores using politeness package in R. We utilized SpaCy (Honnibal et al., 2017) to get number of named entities and root clauses. For discourse connectives, stance adverbials, reasoning verbs, and modals, we adapted the code from Kolhatkar et al. (2020).
After calculating the features, we performed Mann-Whitney tests with Bonferroni correction to inspect whether these features differed between the human-written baseline and constructive comments. However, we found no significant difference (see Table 1). To complement this analysis, we used LLM-as-a-judge (GPT-4) to compare which comment appeared more constructive in a pair, where each pair consisted of one randomly selected comment from the baseline and another from the constructive condition, matched on country, topic, and stance. Across 200 random pairs, the LLM showed no clear preference, selecting the baseline comment as more constructive in 48% of cases and the one from constructive condition in 52% of cases. These findings indicate that even when participants were asked to write constructively, their comments were largely indistinguishable from the baseline comments based on both linguistic features and perceived constructiveness. Therefore, for subsequent analyses we merged both sets and treated them as human-written comments on value-laden topics.
Table 1. Constructive features in human-written comments across baseline and constructive conditions.
Constructive
Feature
Baseline Condition
Constructive Condition
Length 73.3 71.4
Readability 8.8 9.0
Named entity 2.73 2.03
Discourse
connectives
7.28 6.83
Stance
adverbials
0.74 0.69
Reasoning verbs
and modals
5.18 5.15
Root clauses 4.03 4.16
Politeness 13.97 14.91
@@
Value Analysis. To examine how participants expressed their values, we annotated the underlying values in different segments of each comment, following Schwartz’s theory of basic human values (Schwartz, 1994). Schwartz’s theory has been used to detect values in arguments in cross-cultural setting (Kiesel et al., 2022). Additionally, researchers applied this framework to measure human-LLM value alignment (Shen et al., 2024, 2025; Liu et al., 2025; Wang et al., 2025; Segerer, 2025). Schwartz’s theory identifies a set of universal values that guide human behavior, organized into ten broad categories such as Security, Tradition, Conformity, Universalism, Benevolence, Power, Achievement, Stimulation, Hedonism, and Self-direction (Schwartz, 1994) (see Figure 3). These values are grouped under high level categories, such as Openness to change (open to new views) vs. Conservation (opposing change) and Self-enhancement (pursuing own interest) vs. Self-transcendence (caring for others).
Figure 3. Circular motivational continuum of Schwartz’s basic human values (adapted from (Schwartz et al., 2012)).
A circular visualization of Schwartz’s basic human values. Openness to change: Self-direction, Stimulation, Hedonism, Self-enhancement: Power, Achievement, Hedonism, Conservation: Face, Security, Tradition, Conformity, Self-transcendence: Universalism, Benevolence, Humility
First, we tested with GPT-4, Gemini, and Claude to annotate a small sample of comments. We noticed that Gemini frequently applied labels outside Schwartz’s framework, while GPT-4 and Claude produced more consistent annotations. We selected GPT-4 as the reference model because prior research shows that it achieves higher accuracy than both Claude (Nemkova et al., 2025) and expert crowdworkers (He et al., 2024) in diverse annotation tasks. We used the following prompt:
Identify Schwartz’s human values in different segments of the following comment.
¡insert comment¿
Only list the segments and corresponding value pairs in the following format. Don’t add any explanation.
Segment:
Value:
@@
Following the initial annotation process, each comment represented a vector of values. For example, the comment shown in Figure 4 represents six distinct values: Universalism, Hedonism, Self-direction, Tradition, Conformity, and Security. The annotations can either have a positive or negative association with the value. For example, the segment ‘why society should interfere’ captures lack of tolerance (Universalism). Some segments can represent multiple values, such as ‘man get married to a woman’ signifies both Traditional value and Conformity to rules.
@@
To verify whether GPT-4 correctly recognized values across different cultural contexts, two of the authors independently reviewed all the annotations. The authors bring socio-cultural expertise from their longstanding research experience in India and lived experience in the US. We revised GPT-4’s annotations when we disagreed with the assigned value or when a comment segment reflected multiple values that GPT-4 did not capture. Inter-rater reliability between authors was high (Cohen’s $\kappa=0.87$) and remaining disagreements were resolved through discussion. After confirming the annotations, we conducted multiple CHI-square tests with Bonferroni correction to examine cross-cultural differences in the values expressed by Indian and American participants across different topics and stances. We also examined whether the values differed based on different demographic characteristics.
@@ 3.2. Phase 2: Values in LLM-Rewritten Constructive Comments
LLM Rewriting. We used GPT-4 to rewrite participants’ comments from Phase 1 constructively. To minimize misalignment between human’s and LLM’s understanding of constructiveness, we relied on different characteristics of constructive comments that participants provided in the constructive condition of Phase 1. We synthesized their open-ended responses into a summary definition: constructive comments are “respectful, fact-based, balanced, and thoughtful contributions that promote understanding and progress in conversation rather than conflict”–which also align with the perceptions of constructiveness from prior studies (Sukumaran et al., 2011; Friess and Eilders, 2015; Kolhatkar and Taboada, 2017a; Napoles et al., 2017; Zhang et al., 2018). We used zero-shot, cultural prompting to ensure that LLM rewrites algin with the corresponding country’s cultural values (Tao et al., 2024). We instructed GPT-4 to rewrite each comment constructively within 100 words, matching the average length of constructive comments as reported in prior studies (Kolhatkar and Taboada, 2017b). We used the following prompt:
@@
For each human-written comment from Phase 1, we generated one LLM-rewritten version, resulting in 234 LLM-rewritten constructive comments. Since we used a fixed prompt to generate these rewrites, we wanted to make sure that our findings are not sensitive to the prompt. Therefore, we collected additional comments from participants, where they could choose their own prompts to rewrite constructively using LLM.
Human-LLM Collaborative Writing. Similar to Phase 1, we recruited 51 new participants (US: 26, India: 25) on Prolific for writing comments on value-laden topics. Participants first reported their stance on same-sex marriage and Islam since people’s values and stances on value-laden topics are closely related (Mason, 2015). Then they were asked to write a comment in response to a randomly selected homophobic and an Islamophobic thread. We used different sets of threads (see Table 10 in Appendix) from Phase 1 to ensure that our analyses were not sensitive to specific threads.
@@
We asked participants to write a comment on their own first. Then they could prompt an LLM to rewrite their comments constructively (see Figure 5). We provided a set of prompts based on the characteristics of constructive comments that other participants described in Phase 1. Participants could select multiple prompts or write their own prompt. We used GPT-4 to rewrite their comments constructively in real-time using the following instruction:
@@
Data Analysis. We generated 234 LLM-rewritten version of the comments that participants wrote in Phase 1. Additionally, from human-LLM collaborative writing in Phase 2, we collected 102 human-written comments along with LLM-rewritten versions submitted by participants. Overall, this led to 336 pairs of human-written comments and their corresponding LLM-rewritten versions.
First, we wanted to verify if the LLM-rewritten versions were indeed more constructive than the original human-written comments. For this, we calculated the linguistic features of constructiveness in comments as described in Phase 1 and performed Mann-Whitney tests with Bonferroni correction. Second, to examine how LLMs handle underlying values when rewriting constructively, we annotated the values in comments based on Schwartz’s theory of basic human values (Schwartz, 1994). We followed the same procedure as described above in Phase 1. We performed multiple CHI-square tests with Bonferroni correction to analyze whether the values differed between human-written and LLM-rewritten constructive comments.